OpenStudy (anonymous):

The center of a circle is (4, 6), and an endpoint of a diameter is (2, 5). What is the other endpoint of the diameter? a. (0, 4) b. (3, 5.5) c. (6, 7)

OpenStudy (anonymous):

the distance between the center and an endpoint of a diameter is the radius. So, just get the distance from center to the endpoint of a diameter then multiply it by 2

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

let (4,6) be one endpoint and (2,5) another endpoint. use distance formula. \[d = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\]

OpenStudy (anonymous):

by d i mean distance not diameter

OpenStudy (anonymous):

to find the diameter just multiply it by 2 \[diameter = 2\sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\]

OpenStudy (anonymous):

i got \[\sqrt{5}\]

OpenStudy (anonymous):

that's right. but that's just the distance from the center to one of the endpoint of the diameter. like i said, you have to multiply it by 2 to get the length of the diameter.

OpenStudy (anonymous):

\[\sqrt{10} = \sqrt{5}\sqrt{2}?\]

OpenStudy (anonymous):

multiply by 2. not square root 2.

OpenStudy (anonymous):

yeah \[\sqrt{10}???? \sqrt{5}\sqrt{2}\] is just in simplified form? :S

OpenStudy (anonymous):

you multiplied square root of 5 by square root of 2. you just multiply square root of 5 by 2 not square root of 2

OpenStudy (anonymous):

so \[2 \sqrt{5}\]???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but arent there 2 coordinates to find?

OpenStudy (anonymous):

i forgot it was looking for a coordinate sorry.

OpenStudy (anonymous):

so its not like this?

OpenStudy (anonymous):

it's close....okay..it's not like that sorry...this uses midpoints instead

OpenStudy (anonymous):

the center is the midpoint of the diameter. so let's say the other endpoint of the diameter is (x,y). (4,6) is the midpoint of the segment from (x,y) to (2,5) you know the formula for midpoint right? \[m_x = \frac{x_1 + x_2}{2}\] \[m_y = \frac{y_1 + y_2}{2}\]

OpenStudy (anonymous):

so mx = 4 my = 6 x1 = x y1 = y x2 = 2 y2 = 5 now substitute

OpenStudy (anonymous):

sorry for my earlier mistake

OpenStudy (anonymous):

no worries :)

OpenStudy (anonymous):